A small plane is flying a banner in the shape of a rectangle.the area of the banner is 144 square feet.the width of the banner 1/4 the length of the banner.what are the dimensions of the banner.
Answer by stanbon(75887)

A small plane is flying a banner in the shape of a rectangle.the area of the banner is 144 square feet.the width of the banner 1/4 the length of the banner.what are the dimensions of the banner.
--------------------
Area = length*width
-----
144 = L*(L/4)
-----
4*144 = L^2
----
Length = 2*12 = 24'
------
Width = (1/4)24 = 6'
-------------------
Cheers,
Stan H.
-----------------
-----